OpenStudy (anonymous):

What value of m makes the equation true? [–6 + (–13)] + 8 = –6 + (m + 8) m =

OpenStudy (anonymous):

@jennyrlz

OpenStudy (kendricklamar2014):

What do you think the answer is?

OpenStudy (anonymous):

um -8 idk

OpenStudy (kendricklamar2014):

Incorret, I will show you the steps :)

OpenStudy (anonymous):

Do you know what PEMDAS is??

OpenStudy (anonymous):

a wt and no

OpenStudy (anonymous):

ok @KendrickLamar2014 lets do this

OpenStudy (kendricklamar2014):

OK.....

OpenStudy (anonymous):

PEMDAS is the order of operations in order to complete this. Parentheses are solved first Exponets (There are none in this) Multiplication/Division - Whichever come first in the equation Addition/Subtraction - Same as above

OpenStudy (anonymous):

That is how you are gonna need to complete this

OpenStudy (anonymous):

oh yea oops i forgot lol

OpenStudy (kendricklamar2014):

1. We simplify the equation: -6 −13 + 8 = −6 + m + 8 Which simplify's too: −6 + − 13 + 8 = −6 + m + 8

OpenStudy (kendricklamar2014):

2. We combine like terms. Which gives you: (−6+−13+8)=(m)+(−6+8)

OpenStudy (kendricklamar2014):

Which equals: \[−11=m+2\]

OpenStudy (kendricklamar2014):

3. We flip the equation to make it look like:\[m+2=−11\]

OpenStudy (kendricklamar2014):

4. We subtract 2 from both sides:\[m+2−2=−11−2\] Which equals:\[?\]

OpenStudy (jennyrlz):

i am here now :)

OpenStudy (kendricklamar2014):

What do that equal @WOLFPACK13

OpenStudy (jennyrlz):

oh it seems like you are already being helped :)

OpenStudy (jennyrlz):

good luck :)